OpenStudy (anonymous):

which is the equation of the line that passes through (-2;-1) and (-4;-3)

OpenStudy (anonymous):

Find the slope, then use the point/slope form to get the equation of the line. If you need it in slope/intercept form or general form, you will have to do a little algebra to make it fit the desired specification.

OpenStudy (anonymous):

Is that enough information for you to get this one?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Find the slope:\[m=\frac{y_2-y_1}{x_2-x_1}=\frac{-3+1}{-4+2}=1\]

OpenStudy (anonymous):

Point/slope form\[y-y_1=m(x-x_1) \implies y+3=1(x+4) \implies y=x+1 \implies x-y=-1\]

OpenStudy (anonymous):

The equations given there are the line in point/slope form, slope/intercept form, and general form.

OpenStudy (anonymous):

How is that?